Geometry. If the sides of a square are decreased by 2cmm the area is
decreased by 36cm^2. What were the dimensions of the original square?''.

Answer by ptaylor(1894)

Geometry. If the sides of a square are decreased by 2cmm the area is
decreased by 36cm^2. What were the dimensions of the original square?''.


Let x=length of each side of the original square
x^2= area of original square
Then x-2 =length of each side of the smaller square
(x-2)^2= area of smaller square
Now we are told that x^2-(x-2)^2=36 cm^2 clear parens
x^2-x^2+4x-4=36 add 4 from both sides
4x=40
x=10 cm------------------------------length of sides of original square
CK
10^2-8^2=36
100-64=36
36=36


Hope this helps-----ptaylor